# SLOPE and Designing Robust Studies for Generalization

### v1

@@ E.6 Estimating SLOPE for The Median
$\widehat{\textnormal{SLOPE}}^{\textnormal{W}}\left(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{med}}}\right)=$ $\dfrac{\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\left\{\widehat{F}_{P_{O\mid X}}(\widehat{m}_{1/2}\mid X_{i})\widehat{\mu}(X_{i})-O_{i}\mathds{1}(O_{i}\leq\widehat{m}_{1/2})\right\}}{\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\widehat{f}_{P_{O\mid X}}(\widehat{m}_{1/2})},\text{ and}$
$\widehat{\textnormal{SLOPE}}^{\textnormal{R}}\left(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{med}}}\right)=$ $\dfrac{\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\widehat{F}_{P_{O\mid X}}(\widehat{m}_{1/2}\mid X_{i})\widehat{\mu}(X_{i})-\widehat{\textnormal{E}}_{P_{O\mid X}}\left\{O_{i}\mathds{1}(O_{i}\leq\widehat{m}_{1/2})\mid X_{i}\right\}}{\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\widehat{f}_{P_{O\mid X}}(\widehat{m}_{1/2})},$
@@ E.7 Example Estimating Equations for $\omega(x)$
In this section, we discuss some common approaches to estimating the density ratio $\omega(x)$, including the logistic regression, entropy balancing, and a method based on empirical distributions when the support of $Q_{X}$ (i.e., $\mathcal{S}_{X}$) is finite and discrete. Specifically, following the notation in Section E.2, we build estimating equations $g_{5}^{\textnormal{W}}$ for the nuisance parameter $\eta_{5}$ and after obtaining the estimate $\widehat{\eta}_{5}$, we present $\widehat{\omega}(x)$ in terms of $\widehat{\eta}_{5}$. For the sake of exposition, we take the convention that $X$ does not include one (i.e., a constant that can serve as an intercept in regression models).
@@ F.1 Causal Identification
On the source population, we assume SUTVA holds (Assumption 4). Let $X$ be the baseline measurement and $V$ be the village that a household belongs to. We assume the strong ignorability assumption holds with $X$ and $V$, i.e., Assumption 3. We note that this holds by the design of Banerjee et al., (2015). Under these identification assumptions, the SLOPE for the mean becomes
@@ F.2 Estimation of The SLOPE
where $v$ ranges in all villages in the source country and $x\in\mathcal{S}_{X}$ ranges in the categories of the baseline variable, and regression coefficients are constrained to guarantee identification: $\beta_{1}=0$, $\delta_{0}=0$, $(\beta\delta)_{x0}=0$ for $x\in\mathcal{S}_{X}$ and $(\beta\delta)_{0a}=0$ for $a\in\{0,1\}$. In addition, we assume that the outcome variance depends on the baseline measurement $x$ as below.
@@ F.3.1 Auxiliary Data Information
Figure 6 provides the normal diagnostics for the conditional normal assumption imposed in Section 5.2 when estimating the SLOPE for the median. From these QQ-plots, the normality assumption is reasonable well for most countries, but less than perfect for India because of the tail behavior.
@@ H.2 Proof of Theorem 3
$\phi^{\prime}_{\gamma}(0)=\textnormal{W}{Q}_{O,X}\in l^{\infty}(\mathcal{S}_{O,X}),\text{ such that }\int_{B}d\textnormal{W}{Q}_{O,X}=\int_{B}\{O-\textnormal{E}_{P_{O\mid X}}(O\mid X)\}dQ_{O,X}^{0},$
@@
$=\int_{B}\left[\dfrac{\exp(th)}{\textnormal{E}\left\{\exp(thO)\mid X\right\}}\right]dQ_{O,X}^{0}.$
@@
$=\left\{1+thO+O(t^{2})\right\}\left\{1-th\textnormal{E}_{P_{O\mid X}}(O\mid X)+O(t^{2})\right\}-1$
@@
H.3 Proof of Lemma 1
@@
$\psi^{\xi}(Q_{O,X}^{\gamma})=\textnormal{E}_{Q_{X}}\left[\dfrac{\textnormal{E}_{P_{O\mid X}}\left\{\exp(\gamma O)\xi(O,X)\right\}}{\textnormal{E}_{P_{O\mid X}}\left\{\exp(\gamma O)\right\}}\right].$
@@
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\xi})$ $=\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{O\mid X}}\left\{O\xi(O,X)\right\}-\textnormal{E}_{P_{O\mid X}}\left\{\xi(O,X)\right\}-\textnormal{E}_{P_{O\mid X}}(O)\right]$
$=\textnormal{E}_{Q_{X}}\left\{\textnormal{Cov}_{P_{O\mid X}}\left[O,\xi(O,X)\right]\right\}.$
@@ H.4 Proof of Theorem 5
$=\textnormal{SLOPE}(Q_{O,X}^{0},\psi)\cdot\int f_{P_{O\mid X}}(m_{q})dQ_{X}+\textnormal{E}_{Q_{O}^{0}}\{O\mathds{1}(O\leq m_{q}\}-\textnormal{E}_{Q_{X}}\left\{F_{P_{O\mid X}}(m_{q})\mu(X)\right\}$
@@ H.7 Proof of Theorem 8
$=$ $\left\{\begin{bmatrix}1&\textnormal{E}_{Q_{A}}(A)&\textnormal{E}_{Q_{Z}}(Z^{\intercal})\\ \textnormal{E}_{Q_{A}}(A)&\textnormal{E}_{Q_{A}}(A^{2})&\textnormal{E}_{Q_{X}}(AZ^{\intercal})\\ \textnormal{E}_{Q_{Z}}(Z)&\textnormal{E}_{Q_{X}}(ZA)&\textnormal{E}_{Q_{Z}}(ZZ^{\intercal})\end{bmatrix}\right\}^{-1}\begin{bmatrix}&\textnormal{E}_{Q_{X}}\{\sigma^{2}(X)\}\\ &\textnormal{E}_{Q_{X}}\{A\sigma^{2}(X)\}\\ &\textnormal{E}_{Q_{X}}\{Z\sigma^{2}(X)\}\end{bmatrix}$
@@
where entries as “” are omitted because that matrix is symmetric, $b=[\textnormal{E}_{Q_{A}}(A),\textnormal{E}_{Q_{L}}(L^{\intercal})]$, $\delta=\textnormal{Cov}_{Q_{L,A}}[L,A]/\textnormal{Var}_{Q_{A}}[A]$, $V=\textnormal{Cov}_{Q_{X}}(X)-\delta\delta^{\intercal}\textnormal{Var}_{Q_{A}}(A)$, and $S=\begin{bmatrix}\textnormal{Var}_{Q_{A}}(A)&\textnormal{Cov}_{Q_{X}}[A,L^{\intercal}]\\ \textnormal{Cov}_{Q_{X}}[L,A]&\textnormal{Cov}_{Q_{L}}(L)\end{bmatrix}$. Then $\textnormal{SLOPE}(Q_{O,X}^{0},\tau_{\textnormal{a}})$ in (23) follows by expanding terms in the formula above and taking the second entry.
@@ H.8 Proof of Theorem 9
$0=$ $\textnormal{E}_{Q_{X}}\left[f_{P_{O\mid X}}\left(m_{1/2}+{\textnormal{MAD}}\right)\left\{\dfrac{\partial\left\{\psi^{{\textnormal{med}}}(Q_{O}^{\gamma})+\psi^{{\textnormal{MAD}}}(Q_{O}^{\gamma})\right\}}{\partial\gamma}\Bigg|_{\gamma=0}\right\}\right]$
$-\textnormal{E}_{Q_{X}}\left[f_{P_{O\mid X}}\left(m_{1/2}-{\textnormal{MAD}}\right)\left\{\dfrac{\partial\left\{\psi^{{\textnormal{med}}}(Q_{O}^{\gamma})-\psi^{{\textnormal{MAD}}}(Q_{O}^{\gamma})\right\}}{\partial\gamma}\Bigg|_{\gamma=0}\right\}\right]$
@@
H.9 Proof of Theorem 10
Proof of Theorem 10. For the L-functional
@@ H.9 Proof of Theorem 10
We note that SLOPE of L-functionals can be alternatively derived from the IF below (Huber,, 1981, (3.11)) using Theorem 3,
@@ I.1 Proof of Theorem 11 and Theorem 12
The proof for Theorems 11 and 12 follows from standard M-estimation theory in Newey and McFadden, (1994) and Van der Vaart, (2000). We next prove Theorem 11 and omit the proof of Theorem 12 since the proof for both theorems follows the same procedure.
Proof of Theorem 11. We begin by proving consistency under Condition 7 using the standard techniques in Newey and McFadden, (1994, Theorem 2.1). Let
$M(\eta)$ $=-E\left[G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)\right]^{\intercal}E\left[G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)\right],\text{ and}$
$\widehat{M}_{n}(\eta)$ $=-\left\{\dfrac{1}{n}G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)\right\}^{\intercal}\left\{\dfrac{1}{n}G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)\right\}.$
First, under Condition 7(i), $\eta^{\textnormal{W}}$ uniquely maximizes $M(\eta)$. Next, using Lemma 2.4 of Newey and McFadden, (1994), under Condition 7(ii) and (iii), we have the uniform convergence of $G^{\textnormal{W}}$:
$\sup_{\eta\in\Theta}\left\lVert\dfrac{1}{n}\sum_{i=1}^{n}G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)-\textnormal{E}\{G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)\}\right\rVert\to_{p}0,$
and that the function $\textnormal{E}[G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)]$ is continuous with respect to $\eta$, where $\to_{p}$ denotes convergence in probability. Then $M(\eta)$ is also continuous with respect to $\eta$. Next we show the uniform convergence of $\widehat{M}_{n}$. Note that the compactness of $\Theta$ implies the boundedness of $\textnormal{E}[G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)]$. Then for any $\eta\in\Theta$,
@@
$\leq$ $\left\lVert\textnormal{E}\left[G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)\right]-\dfrac{1}{n}\sum_{i=1}^{n}G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)\right\rVert\cdot\ \left\lVert\textnormal{E}\left[G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)\right]+\dfrac{1}{n}\sum_{i=1}^{n}G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta)\right\rVert$
@@
$M(\widehat{\eta}^{\textnormal{W}})>\widehat{M}_{n}(\widehat{\eta}^{\textnormal{W}})-\varepsilon/3>\widehat{M}_{n}(\eta^{\textnormal{W}})-\varepsilon/3-\varepsilon/3>M(\eta^{\textnormal{W}})-\varepsilon/3-\varepsilon/3-\varepsilon/3,=M(\eta^{\textnormal{W}})-\varepsilon.$ (57)
where the first inequality and third equality hold by the uniform convergence (56), and the second inequality holds because $\widehat{\eta}^{\textnormal{W}}$ uniquely maximizes $\widehat{M}_{n}(\cdot)$. Let $\mathcal{N}$ be an open subset of $\Theta$ that contains $\eta^{\textnormal{W}}$. By the compactness of $\Theta$ and continuity of $M$, we have $\sup_{\eta\in\Theta\cap\mathcal{N}^{C}}M(\eta)<M(\eta^{\textnormal{W}})$ with probability approaching one. Let $\varepsilon=M(\eta^{\textnormal{W}})-\sup_{\eta\in\Theta\cap\mathcal{N}^{C}}M(\eta)$, then (57) implies that $\widehat{\eta}^{\textnormal{W}}\in\mathcal{N}$ with probability approaching one. Hence, with Condition 8(iv), the consistency of the SLOPE follows from the continuous mapping theorem.
Next, we prove the asymptotic normality. It directly follows from Theorem 5.31 of Van der Vaart, (2000) that under Conditions 7-8, $\widehat{\eta}^{\textnormal{W}}$ is asymptotically normal in that $\sqrt{n}\left(\widehat{\eta}^{\textnormal{W}}-\eta^{\textnormal{W}}\right)\to_{d}N\left(0,V^{\textnormal{W}}(V^{\textnormal{W}})^{\intercal}\right)$. Then the asymptotic normality of the SLOPE estimate follows from delta method (Theorem 3.1 of Van der Vaart, (2000)).

### v2

@@ E.6 Estimating SLOPE for the Median
$\widehat{\textnormal{SLOPE}}^{\textnormal{W}}\left(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{med}}}\right)=$ $\dfrac{\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\left\{\widehat{F}_{P_{O\mid X}}(\widehat{m}_{1/2}\mid X_{i})\widehat{\mu}(X_{i})-O_{i}\mathds{1}(O_{i}\leq\widehat{m}_{1/2})\right\}}{\dfrac{1}{n_{p}}\sum_{i=n_{q}+1}^{n}\widehat{\omega}(X_{i})\widehat{f}_{P_{O\mid X}}(\widehat{m}_{1/2}\mid X_{i})},\text{ and}$
$\widehat{\textnormal{SLOPE}}^{\textnormal{R}}\left(\widehat{Q}_{O,X}^{0},\widehat{\psi}^{{\textnormal{med}}}\right)=$ $\dfrac{\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\widehat{F}_{P_{O\mid X}}(\widehat{m}_{1/2}\mid X_{i})\widehat{\mu}(X_{i})-\widehat{\textnormal{E}}_{P_{O\mid X}}\left\{O_{i}\mathds{1}(O_{i}\leq\widehat{m}_{1/2})\mid X_{i}\right\}}{\dfrac{1}{n_{q}}\sum_{i=1}^{n_{q}}\widehat{f}_{P_{O\mid X}}(\widehat{m}_{1/2}\mid X_{i})},$
@@ E.7 Example Estimating Equations for $\omega(x)$
In this section, we discuss some common approaches to estimating the density ratio $\omega(x)$, including the logistic regression, entropy balancing, and a method based on empirical distributions when the support of $Q_{X}$ (i.e., $\mathcal{S}_{X}$) is finite and discrete. Specifically, following the notation in Section E.2, we build estimating equations $g_{5}^{W}$ for the nuisance parameter $\eta_{5}$ and after obtaining the estimate $\widehat{\eta}_{5}$, we present $\widehat{\omega}(x)$ in terms of $\widehat{\eta}_{5}$. For the sake of exposition, we take the convention that $X$ does not include one (i.e., a constant that can serve as an intercept in regression models).
@@ F.1 Causal Identification
On the source population, we assume SUTVA (Assumption 4) holds for potential outcome $Y(a)$ under treatment ($a=1$) and control ($a=0$). Let $X$ be the baseline measurement and $V$ be the village that a household belongs to. We assume the strong ignorability assumption holds with $X$ and $V$, i.e., Assumption 3; note that this holds by the design of Banerjee et al., (2015). Under these identification assumptions, the SLOPE for the mean becomes
@@ F.2 Estimation of the SLOPE
where $v$ is a discrete variable ranges in all villages in the source country and $x\in\mathcal{S}_{X}$ ranges in the categories of the baseline variable, and regression coefficients are constrained to guarantee identification: $\beta_{1}=0$, $\delta_{0}=0$, $(\beta\delta)_{x0}=0$ for $x\in\mathcal{S}_{X}$ and $(\beta\delta)_{0a}=0$ for $a\in\{0,1\}$. In addition, we assume that the outcome variance depends on the baseline measurement $x$ as below.
@@ F.3.1 Auxiliary Data Information
Figure 6 provides the normal diagnostics for the conditional normal assumption imposed in Section 5.2 when estimating the SLOPE for the median. From these QQ-plots, the normality assumption is reasonable well for most countries.
@@ H.2 Proof of Theorem 3
$\phi^{\prime}_{\gamma}(0)=\widetilde{Q}_{O,X}\in l^{\infty}(\mathcal{S}_{O,X}),\text{ such that }\int_{B}d\widetilde{Q}_{O,X}=\int_{B}\{O-\textnormal{E}_{P_{O\mid X}}(O\mid X)\}dQ_{O,X}^{0},$
@@
$=\int_{B}\left[\dfrac{\exp(th)}{\textnormal{E}_{P_{O\mid X}}\left\{\exp(thO)\mid X\right\}}\right]dQ_{O,X}^{0}.$
@@
$=\left\{1+thO+O(t^{2})\right\}\left\{1-th\textnormal{E}_{P_{O\mid X}}(O\mid X)+O(t^{2})\right\}-1+O(t^{2})$
@@
H.3 Proof of Theorem 4
Proof of Theorem 4. The proof follows the same procedure as the proof of Theorem 3 except for replacing the exponential function $\exp(\gamma O)$ with $\rho(O,X,\gamma)$. Specifically, the conclusion of the first step becomes
$\phi^{\prime}_{\gamma}(0)=\widetilde{Q}_{O,X}\in l^{\infty}(\mathcal{S}_{O,X}),\text{ such that }\int_{B}d\widetilde{Q}_{O,X}=\int_{B}\left[\dot{\rho}(O,X,0)-\textnormal{E}_{P_{O\mid X}}\{\dot{\rho}(O,X,0)\mid X)\right]dQ_{O,X}^{0}.$
In the derivation of the first step, we use the Taylor expansion on $\rho(O,X,\gamma)$ around $\gamma=0$,
$dQ_{O,X}^{th}=1+th\dot{\rho}(O,X,0)+O(t^{2}),$
which holds under Condition 7. The rest of the proof follows the same procedure as the proof of Theorem 3.
H.4 Proof of Lemma 1
@@
$\psi^{\xi}(Q_{O,X}^{\gamma})=\textnormal{E}_{Q_{X}}\left[\dfrac{\textnormal{E}_{P_{O\mid X}}\left\{\exp(\gamma O)\xi(O,X)\mid X\right\}}{\textnormal{E}_{P_{O\mid X}}\left\{\exp(\gamma O)\mid X\right\}}\right].$
@@
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\xi})$ $=\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{O\mid X}}\left\{O\xi(O,X)\mid X\right\}-\textnormal{E}_{P_{O\mid X}}\left\{\xi(O,X)\mid X\right\}-\textnormal{E}_{P_{O\mid X}}(O\mid X)\right]$
$=\textnormal{E}_{Q_{X}}\left\{\textnormal{Cov}_{P_{O\mid X}}\left[O,\xi(O,X)\mid X\right]\right\}.$
@@ H.5 Proof of Theorem 6
$=\textnormal{SLOPE}(Q_{O,X}^{0},\psi)\cdot\int f_{P_{O\mid X}}(m_{q}\mid X)dQ_{X}+\textnormal{E}_{Q_{O}^{0}}\{O\mathds{1}(O\leq m_{q}\}-\textnormal{E}_{Q_{X}}\left\{F_{P_{O\mid X}}(m_{q})\mu(X)\right\}$
@@ H.8 Proof of Theorem 9
$=$ $\left\{\begin{bmatrix}1&\textnormal{E}_{Q_{A}}(A)&\textnormal{E}_{Q_{L}}(L^{\intercal})\\ \textnormal{E}_{Q_{A}}(A)&\textnormal{E}_{Q_{A}}(A^{2})&\textnormal{E}_{Q_{X}}(AL^{\intercal})\\ \textnormal{E}_{Q_{L}}(L)&\textnormal{E}_{Q_{X}}(LA)&\textnormal{E}_{Q_{L}}(LL^{\intercal})\end{bmatrix}\right\}^{-1}\begin{bmatrix}&\textnormal{E}_{Q_{X}}\{\sigma^{2}(X)\}\\ &\textnormal{E}_{Q_{X}}\{A\sigma^{2}(X)\}\\ &\textnormal{E}_{Q_{X}}\{L\sigma^{2}(X)\}\end{bmatrix}$
@@
where $b=[\textnormal{E}_{Q_{A}}(A),\textnormal{E}_{Q_{L}}(L^{\intercal})]$, $\delta=\textnormal{Cov}_{Q_{L,A}}[L,A]/\textnormal{Var}_{Q_{A}}[A]$, $V=\textnormal{Cov}_{Q_{X}}(X)-\delta\delta^{\intercal}\textnormal{Var}_{Q_{A}}(A)$, $S=\begin{bmatrix}\textnormal{Var}_{Q_{A}}(A)&\textnormal{Cov}_{Q_{X}}[A,L^{\intercal}]\\ \textnormal{Cov}_{Q_{X}}[L,A]&\textnormal{Cov}_{Q_{L}}(L)\end{bmatrix}$, and entries as “” are omitted because that matrix is symmetric. Then $\textnormal{SLOPE}(Q_{O,X}^{0},\tau_{\textnormal{a}})$ in (24) follows by expanding terms in the formula above and taking the second entry.
@@ H.9 Proof of Theorem 10
$0=$ $\textnormal{E}_{Q_{X}}\left[f_{P_{O\mid X}}\left(m_{1/2}+{\textnormal{MAD}}\mid X\right)\left\{\dfrac{\partial\left\{\psi^{{\textnormal{med}}}(Q_{O}^{\gamma})+\psi^{{\textnormal{MAD}}}(Q_{O}^{\gamma})\right\}}{\partial\gamma}\Bigg|_{\gamma=0}\right\}\right]$
$-\textnormal{E}_{Q_{X}}\left[f_{P_{O\mid X}}\left(m_{1/2}-{\textnormal{MAD}}\mid X\right)\left\{\dfrac{\partial\left\{\psi^{{\textnormal{med}}}(Q_{O}^{\gamma})-\psi^{{\textnormal{MAD}}}(Q_{O}^{\gamma})\right\}}{\partial\gamma}\Bigg|_{\gamma=0}\right\}\right]$
@@
H.10 Proof of Theorem 11
Proof of Theorem 11. For the L-estimand
@@ H.10 Proof of Theorem 11
We note that SLOPE of L-estimands can be alternatively derived from the IF below (Huber,, 1981, (3.11)) using Theorem 3,
@@ I.1 Proof of Theorem 12 and Theorem 13
The proof for Theorems 12 and 13 follows from standard M-estimation theory in Newey and McFadden, (1994) and Van der Vaart, (2000). We next prove Theorem 12 and omit the proof of Theorem 13 since the proof for both theorems follows the same procedure.
Proof of Theorem 12. We begin by proving consistency under Condition 8 using the standard techniques in Newey and McFadden, (1994, Theorem 2.1). Let
$M(\eta)$ $=-E\left[G^{W}(T_{i},O_{i},X_{i},\eta)\right]^{\intercal}E\left[G^{W}(T_{i},O_{i},X_{i},\eta)\right],\text{ and}$
$\widehat{M}_{n}(\eta)$ $=-\left\{\dfrac{1}{n}G^{W}(T_{i},O_{i},X_{i},\eta)\right\}^{\intercal}\left\{\dfrac{1}{n}G^{W}(T_{i},O_{i},X_{i},\eta)\right\}.$
First, under Condition 8(i), $\eta^{W}$ uniquely maximizes $M(\eta)$. Next, using Lemma 2.4 of Newey and McFadden, (1994), under Condition 8(ii) and (iii), we have the uniform convergence of $G^{W}$:
$\sup_{\eta\in\Theta}\left\lVert\dfrac{1}{n}\sum_{i=1}^{n}G^{W}(T_{i},O_{i},X_{i},\eta)-\textnormal{E}\{G^{W}(T_{i},O_{i},X_{i},\eta)\}\right\rVert\to_{p}0,$
and that the function $\textnormal{E}[G^{W}(T_{i},O_{i},X_{i},\eta)]$ is continuous with respect to $\eta$, where $\to_{p}$ denotes convergence in probability. Then $M(\eta)$ is also continuous with respect to $\eta$. Next we show the uniform convergence of $\widehat{M}_{n}$. Note that the compactness of $\Theta$ implies the boundedness of $\textnormal{E}[G^{W}(T_{i},O_{i},X_{i},\eta)]$. Then for any $\eta\in\Theta$,
@@
$\leq$ $\left\lVert\textnormal{E}\left[G^{W}(T_{i},O_{i},X_{i},\eta)\right]-\dfrac{1}{n}\sum_{i=1}^{n}G^{W}(T_{i},O_{i},X_{i},\eta)\right\rVert\cdot\ \left\lVert\textnormal{E}\left[G^{W}(T_{i},O_{i},X_{i},\eta)\right]+\dfrac{1}{n}\sum_{i=1}^{n}G^{W}(T_{i},O_{i},X_{i},\eta)\right\rVert$
@@
$M(\widehat{\eta}^{W})>\widehat{M}_{n}(\widehat{\eta}^{W})-\varepsilon/3>\widehat{M}_{n}(\eta^{W})-\varepsilon/3-\varepsilon/3>M(\eta^{W})-\varepsilon/3-\varepsilon/3-\varepsilon/3,=M(\eta^{W})-\varepsilon.$ (58)
where the first inequality and third equality hold by the uniform convergence (57), and the second inequality holds because $\widehat{\eta}^{W}$ uniquely maximizes $\widehat{M}_{n}(\cdot)$. Let $\mathcal{N}$ be an open subset of $\Theta$ that contains $\eta^{W}$. By the compactness of $\Theta$ and continuity of $M$, we have $\sup_{\eta\in\Theta\cap\mathcal{N}^{C}}M(\eta)<M(\eta^{W})$ with probability approaching one. Let $\varepsilon=M(\eta^{W})-\sup_{\eta\in\Theta\cap\mathcal{N}^{C}}M(\eta)$, then (58) implies that $\widehat{\eta}^{W}\in\mathcal{N}$ with probability approaching one. Hence, with Condition 9(iv), the consistency of the SLOPE follows from the continuous mapping theorem.
Next, we prove the asymptotic normality. It directly follows from Theorem 5.31 of Van der Vaart, (2000) that under Conditions 8-9, $\widehat{\eta}^{W}$ is asymptotically normal in that $\sqrt{n}\left(\widehat{\eta}^{W}-\eta^{W}\right)\to_{d}N\left(0,V^{W}(V^{W})^{\intercal}\right)$. Then the asymptotic normality of the SLOPE estimate follows from delta method (Theorem 3.1 of Van der Vaart, (2000)).